Use the transformation v=52+y(2) + 2 to find the solution to the initial value problem dey (2) = 5 +y (2) +2; y(0) = 5
Answer #1

ah ah 5 you= 51 +4(x)+2; y (0) = Let ua 5x+1000+2 =5+ ata yux) as a y(x) = -5 → 0 By equating @ & is du 5 = 5*+9 (2) +2 dv dx